# Generalization v.s. Memorization: Tracing Language Models’ Capabilities Back to Pretraining Data

### v3

@@ 1 Introduction
To our knowledge, this work presents one of the first comprehensive analyses of LLM capabilities by tracing their origins to pretraining corpora at scale. Our task-gram language model offers a scalable, generalizable approach for understanding LLM behavior across a variety of tasks. [^3] Code will be made available upon acceptance of the paper.
@@ 7 Practical Implications: Prompt Optimization
Memorization 28.7% 3.9%
Generalization 23.2% 4.5%
Table 1: Pythia 12B model accuracy on TriviaQA and GSM8K test set with memorization encouraged task prompt (maximize counts) and generalization encouraged task prompt (minimize counts).
@@
We implement a simple prompt optimizer based on GPT4o and the WIMBD $n$-gram count feedback. More specifically, we instruct GPT4o (Achiam et al., 2023) to rewrite a given task prompt at each iteration, and give the average $n$-gram count in the pretraining corpus of the rewritten prompt to GPT4o in the next iteration as the reward. We instruct GPT4o to maximize this reward if we want to encourage memorization, and instruct it to minimize this reward if we want to encourage generalization. Here, we show a maximization and a minimization result for TriviaQA and GSM8K respectively. We report testing accuracy with the Pythia 12B model in Table 1. The meta prompt we used to perform such optimization and the optimized task prompts is included in Appendix E.
Note that the lengths of the optimized prompts are not significantly different, while TriviaQA significantly benefits from the prompt more similar to the pretraining data, and GSM8K benefits from the prompts less similar. More sophisticated prompt optimization algorithms with more detailed distributional memorization feedback can be designed based on a similar idea. We decide to only include the current simple version of the algorithm in this paper and leave the investigation of other possibilities in future work.
@@ 9 Conclusion
In this paper, we introduce the task-gram language model, a scalable approach for modeling the task-relevant distribution of pretraining data. We trace the capabilities of LLMs back to this data by defining distributional memorization, measured through the Spearman correlation between task-gram LM probabilities and LLM probabilities. Through extensive experiments using Pythia models across four distinct tasks, we find that LLMs tend to memorize more when engaged in simpler, knowledge-intensive tasks, while they generalize more in harder, reasoning-intensive tasks. Our analysis provides a comprehensive examination of the origins of LLM capabilities and offers a scalable framework for investigating the fine-grained task-relevant characteristics of pretraining corpora.
@@ Appendix E Prompt Optimization Details
Dataset Memorization Generalization
TriviaQA Deliver an exact single word or concise phrase in response to the factual question. (avg. 3-gram count: 557948.5) Formulate a distinctive and concise term or phrase to clearly answer the factual question. (avg. 3-gram count: 1714.9)
@@
=Text in the square brackets are comments that are not involved in the actual prompt.

### v4

@@ 1 Introduction
We also propose a prompt optimization algorithm to illustrate the practical implications of our analysis. Using GPT-4, we rewrite task prompts to either maximize $n$-gram overlaps with the pretraining corpus to encourage memorization for knowledge-intensive tasks or minimize these overlaps to promote generalization for reasoning-intensive tasks. Empirical results across Pythia models of varying sizes and OLMo models on TriviaQA and GSM8K validate the effectiveness of the proposed approach.
To our knowledge, this work presents one of the first comprehensive analyses of LLM capabilities by tracing their origins to pretraining corpora at scale. Our task-gram language model offers a scalable, generalizable approach for understanding LLM behavior across a variety of tasks.
@@ 7 Practical Implications: Prompt Optimization
Memorization Generalization Memorization Generalization
Pythia (6.9B) 17% 9% 2.6% 2.8%
Pythia-Instruct (6.9B) 23.5% 23.2% 6.3% 7.3%
Pythia (12B) 28.7% 23.2% 2.7% 2.8%
OLMo (7B) 36.4% 29.8% 2.5% 3.1%
OLMo-instruct (7B) 29% 10% 6.3% 7.9%
Table 1: Zero-shot accuracy on TriviaQA and GSM8K test set with memorization encouraged task prompt (maximize counts) and generalization encouraged task prompt (minimize counts).
@@
We implement a simple prompt optimizer based on GPT4o and the WIMBD $n$-gram count feedback. More specifically, we instruct GPT4o (Achiam et al., 2023) to rewrite a given task prompt at each iteration, and give the average $n$-gram count in the pretraining corpus of the rewritten prompt to GPT4o in the next iteration as the reward. We instruct GPT4o to maximize this reward if we want to encourage memorization, and instruct it to minimize this reward if we want to encourage generalization. Here, we show a maximization and a minimization result for TriviaQA and GSM8K respectively. We report zero-shot testing accuracy with the Pythia models and OLMo models in Table 1. The meta prompt we used to perform such optimization and the optimized task prompts are included in Appendix E.
Note that the lengths of the optimized prompts are not significantly different, while TriviaQA significantly benefits from the prompts that are more similar to the pretraining data, and GSM8K benefits from the prompts that are less similar. More sophisticated prompt optimization algorithms with more detailed distributional memorization feedback can be designed based on a similar idea. We leave the investigation of other possibilities for future work.
@@ 9 Conclusion
In this paper, we introduce the task-gram language model, a scalable approach for modeling the task-relevant distribution of pretraining data. We trace the capabilities of LLMs back to this data by defining distributional memorization, measured through the Spearman correlation between task-gram LM probabilities and LLM probabilities. Through extensive experiments using Pythia models across four distinct tasks, we find that LLMs tend to memorize more when engaged in simpler, knowledge-intensive tasks, while they generalize more in harder, reasoning-intensive tasks. Our analysis provides a comprehensive examination of the origins of LLM capabilities and offers a scalable framework for investigating the fine-grained task-relevant characteristics of pretraining corpora. We also propose a simple prompt optimization algorithm to demonstrate the practical implication of our analysis.
@@ Appendix E Prompt Optimization Details
Corpus Dataset Memorization Generalization
Pile TriviaQA Deliver an exact single word or concise phrase in response to the factual question. (avg. 3-gram count: 557948.5) Formulate a distinctive and concise term or phrase to clearly answer the factual question. (avg. 3-gram count: 1714.9)
@@
Dolma TriviaQA Provide a single word or concise phrase in response to the following factual question. (Avg. 3-gram count: 5566475.5) Provide a clear and specific word or brief phrase in response to the factual question below. (Avg. 3-gram count: 43436.2)
GSM8K Solve the following math word problem by methodically breaking it down into simple, clear steps to find the correct solution efficiently. (Avg. 3-gram count: 3736735.0) Solve the upcoming math word problem by sequentially explaining each calculation and logical step, ensuring clarity and coherence in your solution. (Avg. 3-gram count: 4563.6)
@@
Text in the square brackets are comments that are not involved in the actual prompt.